Twice the sum of a number and 4 gives 10. What is the number?
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52781)

.
2*(x+4) = 10,

2x + 8 = 10,

2x = 10-8 = 2,

x = 1.

Solved.


Answer by josgarithmetic(39617)